# Photonuclear treatment for spent fuel radiotoxicity reduction: a case study investigation on minor actinides

### v1

@@ 4.1 Atomic concentrations evolution
From the analysis reported in Sec. 3.2, both the photon and neutron fluxes from scenarios A and B shown in Figure 6 do not have appreciable differences, meaning that the photon/neutron buildup within the pellet is not sensible to a an external converter. For this reason, from now on just the results concerning scenarios B and C will be shown. The driver fluxes are maintained constant over time, assuming that any feedback from the fuel composition on the photon and neutron transport within the pellet is negligible.
Temporal evolution are shown in Figure 7 and Figure 8 for cases B and C, respectively. Two major behaviours can be appreciated. The first involves those nuclides present in a considerable amount in the initial condition, namely 241Am, 243Am, 244Cm and 237Np (solid lines). They tend to decrease in time, being both transmuted with or reactions or destroyed by fission. Then, the second behaviour involves the nuclides formed through the transmutation of the original isotopes. Several byproducts (indicated by dotted lines) appear in the system, initially starting at low concentrations, peaking when the transmutation effect of the precursors is at its maximum, and then reducing due to the combined effects of transmutation and fission. Among the 53 nuclides studied, only the four primary contributors in terms of mass for each species are presented, as the remaining nuclides are present only in traces.
The shown results consider a photon flux normalized to 1E+18 . Since the flux is a linear parameter in the Bateman equation, the dynamics of nuclide evolution (specifically, how they tend to evolve relative to each others) appears to be unaffected by the power level. Indeed, the consumption of the nuclides is directly proportional to the averaged flux value. For this level, the attenuation of the actinides is achieved in less than one year. A more detailed discussion regarding the quantification of this treatment’s effects will be provided in the next subsection. It’s interesting to notice that similar results in terms of concentration reduction were found by [17] with photon flux targeting fission products.
@@ 4.2 Radiotoxicity estimation
Figure 11 shows the normalized radiotoxicity evolution due to the minor actinide content when the irradiation beam is given by Bremmstrahlung photons (i.e., Simulation B) or by LCB photons (i.e., Simulation C). Both cases reveal a similar trend: the initial radiotoxicity tends to increase up to a factor 5 due to the photon transmutations, leading the system to increase the radiotoxic nuclides content. Then, a rapid decrease is observed due to further transmutations to less radiotoxic species, as well as due to fission disintegrations. Simulation (C) appears to be more effective in terms of timing with respect to Simulation (B), crossing sooner the threshold. The time in which the radiotoxicity curve crosses the value of 1, is called cut-off time. Both curves are compared with the radiotoxicity trend in case of no SNF treatment. It is worth noticing that the evolution in all the three cases starts after two years the fuel was stored in a decay pool.
@@ 5 Conclusions
This paper investigated the methodological approach of adopting photon-guided irradiation to achieve a reduction in minor actinides content, with focus on the physical phenomenon and interaction in the spent fuel. The presented methodology has been applied to spent fuel simulated from a SMR technology, specifically the NuScale design, but it remains general for every reactor concept. The analysis was conducted following a step-by-step approach, taking care to consider only the minor actinides present in the total spent fuel content. Three separate simulations were conducted, each with a different irradiation geometry: electron irradiation using a tantalum converter (A), direct conversion on the actinides (B), and photon irradiation from a laser combined beam (C).
The Bateman equations were solved for 53 nuclides, accounting for their mutual decay chain as well as reactions resulting from interactions with photons and neutrons’ transmutations and fissions. Several volume averaged fluxes have been considered, spanning from 1E+17 to 1E+20 . Through the adoption of damage coefficients for the nuclides involved in the analysis, the ingestion radiotoxicity was calculated and tracked over time, allowing for comparison with natural radiotoxicity. Depending on the flux level employed, the cut-off time for the radiotoxicity of the spent fuel pellet decreases, following a linear trend with power, effectively burning 130 grams of spent fuel in approximately 30 days when a volume averaged flux of 1E+20 is used in simulation (C).
The technologies examined, including LINAC, electron irradiation, and laser-Compton backscattering irradiation, were compared at the same flux level, showing the best performances of the quasi monoenergetic flux shape in the LCB case. However, all methods effectively utilized photonuclear reactions, proving to be viable alternatives for the transmutation of minor actinides. They significantly reduce the time required for these products to decrease to radiotoxicity levels comparable to natural uranium ores, achieving reductions by several orders of magnitude.
While there are opportunities for improvement from a technological perspective, as higher flux intensities correlate with shorter irradiation times, the current analysis suggests that the LCB technology is the most effective in terms of performance, being able to maximize the transmutation capabilities. Overall, these findings underscore the potential of this method in addressing the challenges associated with nuclear waste treatment.

### v2

@@ 4.1 Atomic concentrations evolution
From the analysis reported in Sec. 3.2, both the photon and neutron fluxes from scenarios A and B shown in Figure 6 do not have appreciable differences, meaning that the photon/neutron buildup within the pellet is not sensible to a an external converter. For this reason, from now on just the results concerning scenarios B and C will be shown. The driver fluxes are updated over time with the Forward Euler scheme reported in Algorithm 1: within the considered time frame of 100 years, we performed 5 substeps at [0.0, 0.25], [0.25, 0.50], [0.50, 1], [1, 5] and [5, 100] years where the fluxes are re-evaluated based on the change in composition. The overall nuclides’ evolution in concentration is shown in Figure 7 and Figure 8 for cases B and C, respectively. Two major behaviours can be appreciated. The first involves those nuclides present in a considerable amount in the initial condition, namely 241Am, 243Am, 244Cm and 237Np (solid lines). They tend to decrease in time, being both transmuted with or reactions or destroyed by fission. Then, the second behaviour involves the nuclides formed through the transmutation of the original isotopes. Several byproducts (indicated by dotted lines) appear in the system, initially starting at low concentrations, peaking when the transmutation effect of the precursors is at its maximum, and then reducing due to the combined effects of transmutation and fission. Among the 53 nuclides studied, only the four primary contributors in terms of mass for each species are presented, as the remaining nuclides are present only in traces.
The shown results consider an initial photon flux normalized to 1E+18 . Since the flux is a linear parameter in the Bateman equation, the dynamics of nuclide evolution (specifically, how they tend to evolve relative to each others) appears to be unaffected by the power level. Indeed, the consumption of the nuclides is directly proportional to the averaged flux value. For this level, the attenuation of the actinides is achieved in less than one year. A more detailed discussion regarding the quantification of this treatment’s effects will be provided in the next subsection. It’s interesting to notice that similar results in terms of concentration reduction were found by [18] with photon flux targeting fission products.
@@ 4.2 Radiotoxicity estimation
Figure 11 shows the normalized radiotoxicity evolution due to the minor actinide content when the irradiation beam is given by Bremmstrahlung photons (i.e., Simulation B) or by LCB photons (i.e., Simulation C). Both cases reveal a similar trend: the initial radiotoxicity tends to increase up to a factor 5 due to the photon transmutations, leading the system to increase the radiotoxic nuclides content. Then, a rapid decrease is observed due to further transmutations to less radiotoxic species, as well as due to fission disintegrations. Simulation (C) appears to be more effective in terms of timing with respect to Simulation (B), crossing sooner the threshold. The shown radiotoxicity regards the MAs content only, not considering the transmutations that may occur to LLFPS, eventually increasing the system radiotoxicity.
The time in which the radiotoxicity curve crosses the value of 1, is called cut-off time. Both curves are compared with the radiotoxicity trend in case of no SNF treatment. It is worth noticing that the evolution in all the three cases starts after two years the fuel was stored in a decay pool.
@@ 5 Conclusions
This paper investigated the methodological approach of adopting photon-guided irradiation to achieve a reduction in minor actinides content, with a focus on the physical phenomenon and interaction in the spent fuel. The industrial feasibility (i.e., the cost estimation to maintain the considered fluxes) was not taken into account in this work. The presented methodology has been applied to spent fuel simulated from a SMR technology, specifically the NuScale design, but it remains general for every reactor concept. The analysis was conducted following a step-by-step approach, taking care to consider only the minor actinides present in the total spent fuel content. Three separate simulations were conducted, each with a different irradiation geometry: electron irradiation using a tantalum converter (A), direct conversion on the actinides (B), and photon irradiation from a laser combined beam (C).
The Bateman equations were solved for 53 nuclides with a Forward Euler scheme, accounting for their mutual decay chain as well as reactions resulting from interactions with photons and neutrons’ transmutations and fissions. Several initial volume averaged fluxes have been considered, spanning from 1E+17 to 1E+20 , chosen accordingly to literature reference. Through the adoption of damage coefficients for the nuclides involved in the analysis, the ingestion radiotoxicity was calculated and tracked over time, allowing for comparison with natural uranium radiotoxicity. Depending on the flux level employed, the cut-off time for the radiotoxicity of the spent fuel pellet decreases, following a linear trend with power, effectively burning 130 grams of Minor Actinides present in the spent fuel in approximately 30 days when a volume averaged flux of 1E+20 is used in simulation (C).
The technologies examined, including LINAC, electron irradiation, and laser-Compton backscattering irradiation, were compared at the same flux level, showing the best performances of the quasi monoenergetic flux shape in the LCB case. However, all methods effectively utilized photonuclear reactions, suggesting to potentially become a viable alternatives for the transmutation of minor actinides if such fluxes were maintained in the system. In this scenario, the time required for these products to decrease to radiotoxicity levels comparable to natural uranium is highly reduced by several orders of magnitude.
While there are opportunities for improvement from a technological perspective, as higher flux intensities correlate with shorter irradiation times, the current analysis suggests that the LCB technology would be the most effective in terms of performance, being able to maximize the transmutation capabilities with the considered flux level. Overall, these findings underscore the potential of this method in addressing the challenges associated with nuclear waste treatment.